OpenStudy (swaqqout_kid1):

The blues band, Jonny and the Silver Toads, charges $35 per ticket at their performances. Their next venue charges them $900 for use of the venue. Based on the inequality below, how many tickets, t, do they need to sell in order to profit at least $3,300? $35t - $900 > $3,300 t > 120 t > 69 t > 96 t > 137 @EmmaMink @magepker728

OpenStudy (jdoe0001):

how would you solve say " 35t - 900 = 33000 " ? what would it look like, when solving for "t"?

OpenStudy (anonymous):

The problem to solve is: \[35t - 900 > 3,300\]First, lets work on the left hand side of your inequality, the:\[35t - 900 \]This means, for instance, to see if it can be simplified at all. Multiply: \[t ~ x~35\]The answer would be t. \[35~x~t~evaluates~\to~35t\]So all in all, the left hand side of your inequality can be written as: \[35t - 900 \]Now let's work on the right hand side of your inequality, the 3,300. The right hand side of your inequality can be written as: \[3,300\]So with these simplifications, the inequality we'll set out to solve is: \[35t-900 >3,300\]Move the -900 to the right side by adding 900 to both sides, like this: \[To~The~Left~Hand~Side: -900+900 =0\]The answer is 35t. To the right hand side:\[3,300+900=4,200\]The answer is 4,200. Now the inequality reads: \[35t> 4,200\]The solution to your inequality is: \[t>120\]